OpenStudy (anonymous):

Find the general term for the sequence 6, 9, 12, 15, ....

OpenStudy (anonymous):

+ 3

OpenStudy (anonymous):

You add 3 to each term.

OpenStudy (anonymous):

6+3= 9+3=12+3=15 make sense?

OpenStudy (anonymous):

answer choices : n+ 5 4n+ 2 3n+ 3 6n+ 3

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

plus three isnt a choice. i get what you are saying i thought the same thing...

OpenStudy (anonymous):

It's c 3n + 3

OpenStudy (anonymous):

\[3(1) + 3 = 6\]

OpenStudy (anonymous):

\[3(2)+3 = 9\]

OpenStudy (anonymous):

\[3(3)+3=12\]

OpenStudy (anonymous):

3(4)+3=15 make sense now?

OpenStudy (anonymous):

ahhhh. thank you soo much!! :)

OpenStudy (anonymous):

No problem! :D